Therese is eight years younger than Sian. If the sum of their ages is 34, how old is Tesrese?
Answer by Theo(13342)

let t = age of therese.
let s = age of sian.
t = s - 8
t + s = 34
replace t with s - 8 to get:
s - 8 + s = 34
combine like terms to get:
2s - 8 = 34
add 8 to both sides to get:
2s = 42
solve for s to get:
s = 42/2 = 21
t = s - 8, therefore t = 13.
t = 13
s = 21
t + s = 13 + 21 = 34
s - t = 21 - 13 = 8
you have:
the sum of their ages is 34 and therese is 8 year younger than sian.
solution looks good.
solution is therese is 13 years old.